A neon sign is made of glass tubing whose inside diameter is 2.5 cm and whose length is 5.5 m. If the sign contains neon at a pressure of 1.78 torr at 35°C, how many grams of neon are in the sign? (Hint: Use the volume of cylinder)
1
Expert's answer
2020-11-17T10:33:12-0500

d= 2.5cm

L=5.5m

P=1.78 torr

T=35°c= 308.15K

Volume of cylindrical , "V=2\u03c0r\u00b2L"

V=2 ×π ×(1.25×10-2)2×5.5

V=9.06m3

"PV= nRT"

9.06×1.78 = n× 8.314×308.15

n=0.006moles

Imole =20g

Mass=1.2g